OpenStudy (anonymous):

Simplify: sqrt(18/x^3) and (6x)/(5+sqrt(x))

OpenStudy (anonymous):

\[\sqrt{(18)/(x^3)}\] and \[(6x)/(5+\sqrt{x})\]

OpenStudy (anonymous):

\(\Large \sqrt{18x^3} =\sqrt{(9x^2)(2x)} =3x\sqrt{2x}\) \(\Large \frac{6x}{5+\sqrt{x}} = \frac{6x(5-\sqrt{x})}{(5+\sqrt{x})(5-\sqrt{x})} =\frac{6x(5-\sqrt{x})}{25-x}\)